# Wooly Graphs : A Mathematical Framework for Knitting

### v1

@@ 1 Introduction
Knitting, an art form that weaves loops of yarn into intricate fabrics, has long fascinated both artists and mathematicians. The patterns created through knitting can be highly complex, often challenging to represent and analyze. Recent advances have explored the potential of graphs to model and understand these patterns, offering new perspectives and analytical tools. An example of the knitting, yarn path, and graph can be seen in figure 1. This paper aims to bridge the gap in the literature by discussing methods from graph theory for identifying and categorizing graphs that represent knittable patterns. By defining and analyzing knitting graphs, we can determine whether a given graph corresponds to a knittable pattern. We further explore the properties of these graphs, including Hamiltonian and Eulerian paths, and propose a categorization of knitting techniques based on their graph complexity.
@@
Through this exploration, we hope to provide a framework for understanding and modeling knitting patterns using graph theory, contributing to both the mathematical and artistic communities. In section 1.1 we discuss other crafts that have been modeled by graphs as well as some previous work on knitting. In section 2.1 we define knitting and knitting graphs as well as define some properties 2.2 and define categories of knitting 2.3. In section 3 we discuss how to determine whether a graph is knittable and give some discussion on the uniqueness of a given graph 3.2. We give a discussion of the limitations of this approach in section 4 and conclude in 5.
@@ 1.1 Related Work
Blackwork Embroidery is an embroidery technique that comes from England. It can be modeled as a graph [12] and the paper goes over different properties of Blackwork Embroidery graphs, including defining how the graph must be Eulerian and has parity requirements on the edges.
@@
Hitomezashi, an embroidery technique from Japan, can be encoded using two characters [22]. This paper also discusses the duality property of Hitomezashi, a pattern on one side of the fabric corresponds to the reverse pattern on the other. The authors note that this is also the case in knitting patterns. This is explored with a problem on reversibility in “About Knitting” [1].
@@
Counts models knitting with directed graphs [6] with a focus on visualization using a customized heuristic approach, as well as known graph embedding methods such as multidimensional scaling [7, 23] and the Kamada-Kawaii algorithm [15].
Another way to model knitting is by looking at the individual building blocks in terms of how the yarn is positioned [17]. This paper looks at the topological construction of knitting and breaks it down to the smallest repeatable section. While not about knitting, the book “Figuring Fibers” contains a chapter [5] that discusses a method to understand intermeshed crochet using graph theory. Here we see the graphs are used for understanding how the two meshes fit together. A planar graph that can be crochet will give its dual as locations for the intermeshed portion.
A generative method for creating knitting patterns, KnitGIST[11], takes different design considerations and stitch patterns to create hand-knitting patterns. This work uses a knitting graph, but focuses on the application of creating knitting patterns to fit a given stitch design and shape.
Many works look at machine knitting since it is usually easier to make things at scale with this technique. One paper that starts to bridge the gap [10] begins at hand knitting and converts to machine knitting. This paper also creates methods to combine stitch techniques while keeping the general shape and stretch of the fabric.

### v2

@@ 1 Introduction
Knitting, an art form that weaves loops of yarn into intricate fabrics, has long fascinated both artists and mathematicians. The patterns created through knitting can be highly complex and often challenging to represent and analyze. Recent advances [9, 6] have explored the potential of graphs to model and understand these patterns, offering new perspectives and analytical tools. An example of the knitting, yarn path, and graph can be seen in figure 1. This paper aims to bridge the gap in the literature by discussing methods from graph theory for identifying and categorizing graphs that represent knittable patterns. By defining and analyzing knitting graphs, we can determine whether a given graph corresponds to a knittable pattern. We further explore the properties of these graphs, including Hamiltonian and Eulerian paths, and propose a categorization of knitting techniques based on their graph complexity.
@@
Through this exploration, we hope to provide a framework for understanding and modeling knitting patterns using graph theory, contributing to both the mathematical and artistic communities. In section 1.1, we discuss other crafts that have been modeled by graphs, as well as some previous work on knitting. In section 2.1, we define knitting and knitting graphs as well as define some properties 2.2 and define categories of knitting 2.3. In section 3, we discuss determining whether a graph is knittable and the uniqueness of a given graph 3.2. We give a discussion of the limitations of this approach in section 4 and conclude in 5.
@@ 1.1 Related Work
Blackwork Embroidery is an embroidery technique that comes from England. It can be modeled as a graph [13], and the paper goes over different properties of Blackwork Embroidery graphs, including defining how the graph must be Eulerian and has parity requirements on the edges.
@@
Hitomezashi, an embroidery technique from Japan, can be encoded using two characters [23]. This paper also discusses the duality property of Hitomezashi; a pattern on one side of the fabric corresponds to the reverse pattern on the other. The authors note that this is also the case in knitting patterns. This is explored with a problem on reversibility in “About Knitting” [1].
@@
Counts models knitting with directed graphs [6] with a focus on visualization using a customized heuristic approach, as well as known graph embedding methods such as multidimensional scaling [7, 24] and the Kamada-Kawaii algorithm [16]. This approach has been further developed to better preserve edge lengths and match considerations of knitting geometry [9]
Another way to model knitting is by looking at the individual building blocks in terms of how the yarn is positioned [18]. This paper looks at the topological construction of knitting and breaks it down to the smallest repeatable section. While not about knitting, the book “Figuring Fibers” contains a chapter [5] that discusses a method to understand intermeshed crochet using graph theory. Here, we see the graphs used to understand how the two meshes fit together. A planar graph that can be crochet will give its dual as locations for the intermeshed portion.
A generative method for creating knitting patterns, KnitGIST[12], takes different design considerations and stitch patterns to create hand-knitting patterns. This work uses a knitting graph but focuses on creating knitting patterns to fit a given stitch design and shape.
Many works look at machine knitting since it is usually easier to make things at scale with this technique. One paper that starts to bridge the gap [11] begins with hand knitting and converts to machine knitting. This paper also creates methods to combine stitch techniques while keeping the general shape and stretch of the fabric.